OpenStudy (anonymous):

how do you solve w + 2/5 = 7/10

OpenStudy (anonymous):

subtract 2/5 from each side (which is subtracting 4/10) so w=3/10

OpenStudy (anonymous):

Find a least common denominator, in this case 10. To get a denominator of 10 for 2/5, you have to multiply the top and bottom by 2 to get 4/10, so: w + 4/10 = 7/10 w = 7/10 - 4/10 w = 3/10

OpenStudy (anonymous):

thank you...I get it!